# Pion Valence Structure at Intermediate x in the Residual Field Model

### v1

@@ Abstract
We calculate the valence parton distribution function of the pion, for the range of $0.1\leq x\leq 0.6$, based on the soft dynamics of the partons within the valence-residual field model. Within this model we assume that the soft partonic structure of the pion is defined by the valence current $q\bar{q}$ cluster embedded in the residual field of the pion. The parton distribution function is calculated using phenomenological wave functions for the $q\bar{q}$ cluster and residual field. Our result indicate that the peak position of x weighted valence parton distribution depends on the mass of the residual system and the parameters of the system are obtained by fitting the height and the peak position of this distribution to extracted PDFs. This comparison reveals that unlike for the case of nucleon PDF, very little residual mass is needed to describe existing pion PDFs at intermediate x.
@@ I Introduction
As the lightest hadron and a pseudo-Goldstone boson of chiral symmetry breaking, the pion holds a unique position in hadronic and nuclear physics. Understanding its structure is essential for understanding Quantum Chromodynamics (QCD) nature of strong interaction of nucleons which in part is mediated through the pion exchange. Exploring pion structure at different light-front momentum domain of quarks and gluons will help also to trace the the transition from the partonic to the hadronic degrees of freedom in nuclear interactions at extreme conditions.
One quantity that allows to explore the dynamical structure of pion is its parton distribution function (PDF). PDFs measured over a wide range of x, $10<x<$ 1 can allow to address such questions as the dynamics of the valence distribution as well the role of the sea-quarks and gluons in the pion. It is interesting to explore the role of the valence, sea-quarks and gluons in pion especially from the point of view of their role in mediating interactions between nucleons in nuclei.
Much progress has been made in recent years in the study of pion parton distribution functions (PDFs)Barry et al. (2018a, 2021a). Recent analysis resulted in a high quality valence quark distribution in the pion as well as produced an unexpected result for PDFs at $x\to 1$. At $x\to 1$ if hard gluon exchange plays dominating role in generating PDFs, then within perturbative QCD one predicts an analytic form of pion PDFs to be $(1-x)^{2}$arrar and Jackson (1975); Berger and Brodsky (1979); Gunion et al. (1984). However the mentioned analysis resulted in a behavior of $(1-x)^{\beta}$, with $\beta\approx 1.01$.
@@
Another aspect of the mentioned analysis is obtaining high quality PDF’s at intermediate $x$. For intermediate $x$ (0.1-0.6) the nonperturbative QCD dynamics define the structure of parton PDFs. One prominent feature of valence parton PDFs in this region is that the x weighted partonic distribution ($xf_{i}(x)$) shows prominent peak associated with the average momentum fraction carried by valence quarks. As it was shown in our recent workLeon and Sargsian (2020, 2022) modeling such distribution allows to ascertain the role of the quark-gluon system residual to the valence quarks in the hadron. In Ref.Leon and Sargsian (2020) We developed a model within residual mean field approximation which introduces Light-Front wave functions of valence quark cluster embedded in the residual field of quark-gluon system described by residual LF wave function. Within this model for the case of nucleon it was found that at the starting $Q^{2}_{0}$ when one expects partonic picture to be relevant, the $xf_{i}(x)$ for valence quark PDFs of nucleon peaks at $x_{p}\leq{1\over 4}$ not at ${1\over 3}$ as naively expected. By describing these peaking properties of valence PDFs of nucleons, we found that it requires the residual system to have a mass comparable to pion mass. As such the model indirectly favored the existence of the structure in the nucleon resembling a pion cloud.
In the current work, we develop a residual field model for the description of pion focusing on the domain of $0.1\leq x\leq 0.6$ and investigating the possible properties of the residual system in the pion.
The model divides the hadron into a valence core and a residual structure while utilizing flavor number conservation to treat the valence structure as effectively a system with a fixed number of fermions. The outline for this work is as follows: Sec. II provides a description of the theoretical framework being used and a review of the residual field model. In Sec. III we derive the PDFs based on the model and in Sec. V we compare the numerical estimates of our results with recently obtained phenomenological parameterization of pion PDFs in the range of $0.1\leq x\leq 0.6$. Our focus in this comparison is to evaluate the mass and the distribution of possible residual system in the pion. Sec. VI contains conclusions and outlook. Technical details for certain calculations and specific mathematical definitions are presented in the Appendix.
II Theoretical Framework
In light front (LF) quantization one probes quantum fields at fixed light-cone time, $x^{+}=t+z$, which provides several simplifications in describing high energy processes Brodsky et al. (1998). The LF approach provides a framework to study hadrons in terms of their light front wave functions (LFWFs). In principle, knowing the LFWFs would allow us to specify completely a hadronic state. However, finding the LFWFs would involve solving an infinite set of coupled integral equations, and thus often in practice they are modeled. The modeling of the valence LFWFs for the pion has been done to study the distribution amplitude of the pion Huang et al. (1994). There has also been modeling on the pion based on AdS/CFT correspondence and light front holography Trawiński (2016a, b).
A hadron in a ket state represented by $|P\rangle$ can be described through an expansion in terms of the LFWFs ($\psi_{i/N}\left(\{x_{i},\mathbf{k}_{i}\}_{i}\right)$) Brodsky et al. (1998); Kovchegov and Levin (2012):
$|P\rangle=\sum_{i}\int[d\mu_{i}]\psi_{i/N}\left(\{x_{j},\mathbf{k}_{j}\}_{j}\right)|i\rangle$ (1)
where the sum is over all Fock states $|i\rangle$ and $[d\mu_{i}]$ is the relevant integration measure. Color, spin, etc. degrees of freedom have been suppressed.
One can relate the experimentally measurable PDFs to the LFWFs to constrain them. In the $\overline{MS}$ factorization scheme, the PDF for a quark of flavor $i$ is defined as Sterman et al. (1995); Belitsky and Radyushkin (2005):
$f_{i}(x)=\int\frac{dz^{-}}{2\pi}e^{ixP^{+}z^{-}}\langle P|\bar{\psi}_{i}(z)\gamma^{+}\psi_{i}(0)|P\rangle|_{z^{+},z^{j}_{\perp}=0}$ (2)
Where we have employed the use of the light cone gauge $A^{+}=0$.
Our approach in this work is phenomenological, instead of taking an infinite Fock state expansion as in Eq. 1, we approximate the $|P\rangle$ state through the transition of $N\rightarrow VR$, where $V$ is a valence core and $R$ is a residual state and define phenomenological wave functions in such way that they represent the solution of Weinberg type equation on light-front. Such wave functions are inherently non-perturbative and need to be modeled based on the prominent features of the structure of the hadrons.
@@ II Theoretical Framework
While quark number is not conserved, flavor is. The residual field model utilizes this fact by separating the hadron into a core of valence fermions whose number is conserved and a residual structure which contains everything besides the valence quarks, such as sea quarks and gluons and possibly a meson cloud in the case of the nucleon. This picture is expected to hold for moderate to high x, $x\gtrsim 0.1$. At lower $x$, Regge dynamics dominate and the model likely ceases to be valid. According to residual field model the deep inelastic scattering from a hadron (baryons or mesons) in leading approximation can be described by the Light-Front diagram of Fig.(1(a)) in which scattering takes place with the valence quark from the cluster. The consideration of the residual system in the scattering is equivalent to the spectral function representation of the partonic structure of hadron in which case valence quark distribution indirectly related to the dynamical characteristics of the residual system.
@@ III.1 General Formalism of Deep-Inelastic Scattering
We will consider probing the pion through deep inelastic electron scattering ($e\pi\rightarrow eX$) aimed at probing partonic distribution of valence quarks in the pion. Even though we consider a specific $\pi(e,e^{\prime})X$ reaction the PDFs evaluated in this case are universal and will be the same as PDF’s extracted from Drell-Yan processes in $\pi N\to l^{+}l^{-}+X$ reactions.
Within one-photon exchange approximation the scattering takes place with virtual photon $\gamma^{*}$, with four momenta $q^{\mu}(q_{0},\vec{q})$ where $q_{0}=E_{e}-E^{\prime}_{e}$ and $\vec{q}=\vec{k}-\vec{k}^{\prime}$ with $k^{\mu}(E_{e},\vec{k}_{e})$ and $k^{\mu\prime}(E^{\prime}_{e},\vec{k}^{\prime})$ being four momenta of initial and scattered
The cross section of DIS process in this can be expressed through the contraction of leptonic and hadronic tensors as follows:
${d\sigma\over dQ^{2}dx}={2\pi\alpha^{2}\over Q^{4}}{y^{2}\over Q^{2}}m_{\pi}L_{\mu\nu}W_{\pi}^{\mu\nu}$ (3)
where $Q^{2}=(q^{\mu}q_{\mu})$, $y={p_{\pi}q\over p_{\pi}k_{e}}$ and
$L_{\mu\nu}=4\left[\left(k_{\mu}-{q_{\mu}\over 2}\right)\left(k_{\nu}-{q_{\nu}\over 2}\right)\right]+Q^{2}\left[-g_{\mu\nu}-{q_{\mu}q_{\nu}\over Q2}\right],$ (4)
and the hadronic tensor in general can be defined from deep inelastic electromagnetic current $J^{\mu}(p_{X},s_{X},p_{\pi},s_{\pi})$ in the form:
@@
Further, we will work with light-front momenta for which the 4-vector components are defined as follows:
$k^{\mu}=(k^{+},k^{-},\mathbf{k}^{\perp})=\left(k^{0}+k^{3},k^{0}-k^{3},(k^{1},k^{2})\right)$ (6)
where ‘3’ defines the direction of the large momentum component.
Furthermore we will perform calculations in the Drell-Yan-West reference frame where the 4-momenta of the pion, $p_{\pi}^{\mu}$ , and the virtual photon, $q^{\mu}$, are
$p_{\pi}^{\mu}=\left(p_{\pi}^{+},\frac{m_{\pi}^{2}}{p_{\pi}^{+}},\mathbf{0}^{\perp}\right),\ \ \ q^{\mu}=\left(0,\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}},\mathbf{q}^{\perp}\right).$ (7)
where $Q^{2}=q_{\perp}^{2}$. The high momentum nature of the reaction is reflected in the fact that $p_{\pi}^{+}\gg\sqrt{Q^{2}}\gg m_{\pi}$.
With LF momenta defined in such way the $F_{2}$ structure function of pion is related to the +++++ + component of hadronic tensor as follows:
$F_{2}(x,Q^{2})={m_{\pi}Q^{2}\over 2x(p_{\pi}^{+})^{2}}W_{\pi}^{+,+}$ (8)
and within the leading approximation from the parton model
@@
were $i$ runs for both quarks and anti-quarks in the pion. In the case of the structure function it is dominated by valence quark contribution using the above relations allows to extract sum of quark and anti-quark valence quark distribution weighted by $x$. Usually invoking isospin symmetry it is assumed that the up valence distribution in $\pi^{+}$ is equal to the down valence distribution in $\pi^{-}$: $f_{V}(x,Q^{2})=u^{\pi^{+}}_{V}(x,Q^{2})=d^{\pi^{-}}_{V}(x,Q^{2})$.
III.2 Hadronic Tensor and The Scattering Amplitude
For calculation of pion structure function we consider the diagram of Fig.1(b) in which it is assumed that pion transitions to the cluster of valence $q\bar{q}$ state and residual $R$ state before the interaction of virtual photon with one of the valence quarks in the cluster.
(a) General Hadron
(b) Pion
Figure 1: Feynman Diagram for Deep Inelastic Scattering in Residual-Mean Field model
Assigning to the diagram of Fig. 1(b) an amplitude $A^{\mu}$, for the hadronic tensor, taking into account of the phase spaces of final state quarks and residual system one obtains:
The hadronic tensor, $W^{\mu\nu}$, has the form:

### v2

# Pion Valence Structure at Intermediate x in the Residual Field Approach

@@ Abstract
We calculate the valence parton distribution function (PDF) of pion, within the theoretical approach based on spectral function representation of valence quarks in the pion. In this approach we assume that the soft partonic structure of pion is defined by the valence $q\bar{q}$ cluster, consisting of current quarks, embedded in the residual field of the pion. The valence PDF is calculated using phenomenological Light-Front wave functions for the $q\bar{q}$ cluster and the residual field. Our result indicates that the peak position of the x weighted valence PDF (xPDF) depends on parameters characterizing the virtuality of the cluster and the mass of the residual system. Magnitudes of these parameters are obtained by fitting to the height and the peak position of empirical pion xPDF evaluated at starting $Q_{0}$. They indicate that unlike the nucleon case, very little residual mass is needed to describe existing pion PDFs at intermediate x. They also indicate that the $q\bar{q}$-cluster is highly virtual and on average the interacting quark carries almost all of the momentum of the $q\bar{q}$ cluster. This picture is consistent with the dominance of the Feynman mechanism leaving little room for the hard component in the PDF, and practically describing $\sim(1-x)$ behavior observed recently at $x\to 1$ limit. Our non-trivial observation is that the height and the peak position of xPDF define the analytic behavior of valence PDF at $x\to 1$ limit.
@@ I Introduction
As the lightest hadron and a pseudo-Goldstone boson of chiral symmetry breaking, the pion holds a unique position in hadronic and nuclear physics. Studying its structure is essential for understanding the nature of the strong interaction of nucleons which in part is mediated through the pion exchange.
While in the chiral symmetry breaking picture the long range part of the nucleon-nucleon (NN) interaction is mediated through the near-massless Goldstone field without apparent internal structure of pion, the deep-inelastic scattering from pion indicates its underlying quark-gluon structure which does not reveal itself in NN interaction. This apparent duality indicates on possible specific distribution of valence quarks in pion that makes for example the quark-interchange in NN interaction through the pion field highly unlikely.
One of the important quantities that allows to explore the dynamical structure of pion is its parton distribution function (PDF) probed at the light-front momentum fraction, x ($0<x<1$) and given resolution $Q^{2}$. It allows us to probe valence and sea quark as well as gluon distributions which dominate at different regions of $x$ in the pion. For the valence quarks pion represents the simplest system and as such it is the best testing ground for understanding the QCD mechanism of valence quark interaction in the hadron, thus attracting multitude of theoretical studies (see e.g. Brodsky and Ji (1986); Frederico and Miller (1994); Tiburzi and Miller (2003); Brodsky and de Teramond (2008); Holt and Roberts (2010); Accardi et al. (2010); Pasquini et al. (2023); Ding et al. (2020); Gao et al. (2020)) .
Much progress has been made in the extraction of pion parton distribution exploring variety of deep-inelastic reactions including Drell-Yan, $J/\Psi$- production and Sullivan processes (see. e.g. Owens (1984); Gluck et al. (1992); Sutton et al. (1992); Gluck et al. (1999); Wijesooriya et al. (2005); Aicher et al. (2010); Chekanov et al. (2002); Aaron et al. (2010); Barry et al. (2018a, 2021a, a, 2021a)).
The recent high quality combined analysis of Drell-Yan and Sullivan processesBarry et al. (2018a, 2021a) produced an unexpected result for PDFs at large $x$, observing a behavior of $(1-x)^{\beta}$, with $\beta\approx 1-1.2$. This contradicts to the generally expected dependence of $(1-x)^{2}$ which in part follows from the expectation that high x parton distributions are generated through the hard gluon exchange between valence quarksFarrar and Jackson (1975); Berger and Brodsky (1979); Gunion et al. (1984).
@@
Focusing on the dynamics of valence quark distribution in hadrons, we notice two important properties of valence quarks that will be used in our calculation. The first, is that the valence quarks define quantum numbers of hadrons and their flavor is conserved. As such they can be considered as effective fermions whose number is conserved. The second prominent feature of valence quarks is that their distribution weighted by the momentum fraction $x$, $xf_{i}(x)$ shows a distinctive peak associated with the average momentum fraction carried by valence quarks. In general, such a peaking property in Fermi systems is associated with the “bulk” property of bound system which has an inherently non-perturbative character. In our recent worksLeon and Sargsian (2020, 2022) on modeling valence quark distributions in the nucleon the “bulk” property of hadrons have been described through the quark-gluon system residual to the valence quarks in the nucleon. The model is based on residual mean field approximation which introduces the Light-Front wave function of the valence quark cluster embedded in the interacting residual field. Within this model for the case of the nucleon it was found that at the starting $Q^{2}_{0}=m_{c}^{2}$ ($m_{c}$ is the mass of the charm quark) when one expects the partonic picture to be relevant, the $xf_{i}(x)$ for valence quark PDFs of nucleon peaks at $x_{p}\leq{1\over 4}$ not at $1\over 3$} as naively expected. This observation is confirmed by the modern phenomenological PDFs of nucleon. By describing these peaking properties of the valence PDFs of nucleons, we found that it requires the residual system to have a mass comparable to the pion mass. As such the model indirectly favored the existence of the structure in the nucleon resembling a pion cloud. The model descried reasonably well the valence PDFs for the region $0.2<x<0.6$ underestimating at larger x which indicates the importance of hard mechanism in generation of large x component of valence PDFs in the nucleon.
In the current work we extend the residual field approach for the description of pion valence PDFs. Our goal is to address following main questions: are there structure in pion residual to $q\bar{q}$ valence quarks and how well the mean field distribution of the residual system describes the high x structure of pion PDFs? We introduce new object, that is Light-Front wave function of the residual system which can be investigated further in probing pion form-factors, generatlized partonic and transverse momentum distributions as well as in processes involving target fragmentation in collider kinematics Abir et al. (2023).
The outline for this work is as follows: Sec. II provides the description of the theoretical framework. In Sec. III we derive the PDFs based on the residual field model and in Sec.IV we introduce light-front wave function of two-body system, applying it for $q\bar{q}$-cluster and valence-residual systems respectively. Sec. V presents numerical estimates of the model parameters following from the fit of theoretical expression to the recently obtained phenomenological parameterization of pion PDFs in the range of $0.1\leq x\leq 0.85$. Our focus in this comparison is to evaluate the relevant masses and the light-front momentum distribution of residual system in the pion through the description of peaking property of x weighted pion PDfs at starting $Q_{0}$. We this section we present also the interpretation of the results, Sec. VI contains conclusions and outlook for further application of the presented approach in different processes.
II Theoretical Framework of Residual Field Model
The residual field model utilizes conservation of valence parton quantum numbers to treat them as effective fermions whose numbers are conserved. It separates hadrons into a core of valence fermions consisting of massless quarks and a residual structure which contains everything besides the valence quarks, such as sea quarks and gluons. Interaction with the external probe in this case proceeds through the scattering off the valence quarks in the cluster. This picture is assumed to be valid for moderate to high x, $x\gtrsim 0.1-0.2$ and will break down at lower $x$, where the Regge dynamics dominate in the interaction.
@@ II Theoretical Framework of Residual Field Model
According to the residual field model the deep inelastic scattering from a hadron in the leading approximation can be described by the Light-Front diagram of Fig.(1)(a) in which scattering takes place with a quark inside the valence cluster. The consideration of the residual system in the scattering amplitude is equivalent to the spectral function representation of the partonic structure of the hadron in which case the valence quark distribution is indirectly related to the dynamical characteristics of the residual system.
In light front (LF) quantization one probes quantum fields at fixed light-cone time, $x^{+}=t+z$, which provides several simplifications in describing high energy processes Brodsky et al. (1998). The LF approach provides a framework to study hadrons in terms of their light front wave functions (LFWFs), in which case a hadron in a ket state represented by $|h\rangle$ can be described through the expansion in terms of the LFWFs ($\psi_{i/h}\left(\{x_{i},\mathbf{k}_{i}\}_{i}\right)$) Brodsky et al. (1998); Kovchegov and Levin (2012):
$|h\rangle=\sum_{i}\int[d\mu_{i}]\psi_{i/N}\left(\{x_{j},\mathbf{k}_{j}\}_{j}\right)|i\rangle,$ (1)
where the sum is over all Fock states $|i\rangle$ and $d\mu_{i}]$ is the relevant integration measure. Color and spin degrees of freedom have been suppressed.
With above wave functions the PDF for a quark of flavor $i$, in hadron, $h$ can be represented as the expectaion value of the number operator in the hadronic state Sterman et al. (1995); Belitsky and Radyushkin (2005):
$f_{i}(x,Q^{2})=\int{d^{2}{\bf k_{\perp}}\over(2\pi)^{2}}\langle h(p)\mid b_{i}^{\dagger}(xp,{\bf k_{\perp}})b_{i}(xp,{\bf k_{\perp}})\mid h(p)\rangle,$ (2)
where $b^{\dagger}_{i}$ and $b_{i}$ are creation and annhiliation operators of parton $i$.
Except for the situations of high $Q^{2}$ and $x\to 1$ in which one expects minimal-Fock component to dominate in Eq.(2), the calculation of hadronic PDF’s at moderade $x$ will require LFWFs which are the solution of infinite set of coupled integral equations, containing the different number of quark or gluon constituents. Thus often in practical calculations LFWFs are modeled. The modeling of the valence LFWFs for the pion has been done in different approaches such as Harmonic Oscillator modelsDziembowski and Mankiewicz (1987), models relating equal-time wave function to the light-front wave functionsHuang et al. (1994), light-front consitutent quark modelsPasquini and Schweitzer (2014), the Nambu-Jona Lasinio modelDziembowski and Mankiewicz (1987) as well as model based on light front holography Trawiński (2016a, b).
Our approach in this work is phenomenological, instead of taking an infinite Fock state expansion as in Eq.(1), we approximate the $|h\rangle$ state through the transition of $\pi\rightarrow VR$, where $V$ is a valence $q\bar{q}$ core and $R$ is a residual state that absorbs all higher order Fock-states as well as possible configurations (such as null-mass components) that may couple with $q\bar{q}$-cluster. Furthermore, we define phenomenological wave functions in such a way that they represent the solution of the Weinberg type equation on the light-front for $q\bar{q}$ - cluster state (V), as well as for cluster-residual (VR) effective two body state. Such wave functions are inherently non-perturbative and need to be modeled based on the prominent features of the structure of hadrons.
@@ III.1 General Formalism of Deep-Inelastic Scattering
To set-up derivational framework for pion PDF we consider probing the pion through deep inelastic electron scattering ($e\pi\rightarrow eX$). Using the universality of PDFs the results of calculation can be compared with PDFs extracted for example from Drell-Yan processes in $\pi N\to l^{+}l^{-}+X$ reactions. This is a consequence of the definition of PDFs and validity of QCD factorization for these processes.
Within one-photon exchange approximation the scattering takes place with the virtual photon $\gamma^{*}$, with four momentum $q^{\mu}=k_{e}^{\mu}-k_{e}^{\prime\mu}$, where $k_{e}^{\mu}$ and $k_{e}^{\prime\mu}$ are the four momenta of the initial and scattered electrons respectively.
The cross section for the DIS process can be expressed through the contraction of leptonic and hadronic tensors as follows:
${d\sigma\over dQ^{2}dx}={2\pi\alpha^{2}y^{2}\over Q^{6}}m_{\pi}L_{\mu\nu}W_{\pi}^{\mu\nu},$ (3)
where $Q^{2}=-q^{\mu}q_{\mu}$, $y={p_{\pi}\cdot q\over p_{\pi}\cdot k_{e}}$ and the unpolarized leptonic tensor is
$L_{\mu\nu}=4\left[\left(k_{e,\mu}-{q_{\mu}\over 2}\right)\left(k_{e,\nu}-{q_{\nu}\over 2}\right)\right]+Q^{2}\left[-g_{\mu\nu}-{q_{\mu}q_{\nu}\over Q2}\right].$ (4)
The hadronic tensor, $W_{\pi}^{\mu\nu}$, in general can be defined through the inelastic hadronic current $J^{\mu}(p_{X},s_{X},p_{\pi},s_{\pi})$ in the form:
@@
In derivations, we will work with light-front coordinates with the convention for 4-momenta being:
$k^{\mu}=(k^{+},k^{-},\mathbf{k}^{\perp})=\left(k^{0}+k^{3},k^{0}-k^{3},(k^{1},k^{2})\right).$ (6)
The calculations are performed in the Drell-Yan-West reference frame where the 4-momenta of the pion, $p_{\pi}^{\mu}$ , and virtual photon, $q^{\mu}$ are defined as follows:
$p_{\pi}^{\mu}=\left(p_{\pi}^{+},\frac{m_{\pi}^{2}}{p_{\pi}^{+}},\mathbf{0}^{\perp}\right),\ \ \ q^{\mu}=\left(0,\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}},\mathbf{q}^{\perp}\right),$ (7)
where $q_{\perp}^{2}=Q^{2}$.
In this reference frame the $F_{2}$ structure function of the pion is related to the ++++ component of the hadronic tensor as follows:
$F_{2}(x,Q^{2})={m_{\pi}Q^{2}\over 2x(p_{\pi}^{+})^{2}}W_{\pi}^{+,+},$ (8)
and within the framework of QCD factorization and the definition of PDFs in leading order one can relate the PDF to the structure function $F_{2}$ as:
@@
where $f_{i}(x)$ was defined in Eq.(2) and $i$ runs for quarks or antiquarks in the pion, interacting with the external probe. Comparing Eqs.(8) and (9) one observes that the calculation of PDFs is related to the calculation of the ++++ component of the hadronic tensor, $W_{\pi}^{\mu\nu}$.
In the case of valence quarks using the above relations allows us to extract sum of quark and anti-quark valence quark distributions weighted by $x$. Usually, invoking isospin symmetry it is assumed that the up valence distribution in $\pi^{+}$ is equal to the down valence distribution in $\pi^{-}$: $f_{V}(x,Q^{2})=u^{\pi^{+}}_{V}(x,Q^{2})=d^{\pi^{-}}_{V}(x,Q^{2})$.
III.2 Hadronic Current and Hadronic Tensor
For calculation of pion structure function we consider the diagram of Fig.1(b) in which it is assumed that the pion transitions to the cluster of valence $q\bar{q}$ state and the residual state, $R$ before the interaction of virtual photon with one of the valence quarks in the cluster takes place.
Figure 1: Feynman Diagrams for Deep Inelastic Scattering in Residual-Mean Field model. (a) General case for hadrons with qqqqqq or $q\bar{q}$ clusters. (b) For the case of deep inelastic scattering from pion.
Identifying the hadronic current $J^{\mu}$ in Eq.(5) with the scattering amplitude $A^{\mu}$, which corresponds to the part of the diagram of Fig.1(b) without electron lines and virtual photon propagator and taking into account the phase space of final state quarks and residual system one obtains for the hadronic tensor: